Hello, can you please help me with this question? A ball is thrown upward with an initial speed of 19.6 m/s. When is the ball opposite a roof top that is 14.7 m high? Use the formula h = rt - 4.9t^2 where h is the height.
Answer by ikleyn(52776)

.

To find time "t", you need solve this quadratic equation

    -4.9t^2 + 19.6t = 14.7  meters.


Simplify it step by step

    -4.9t^2 + 19.6t - 14.7 = 0


Divide by (-4.9) both sides

    t^2 - 4t + 3 = 0


Factor left side

    (t-3)*(t-1) = 0.


There are two positive roots  t= 1 second and  t= 3 seconds.


First root is for the ball moving vertically up; the second root is for the ball falling down.

Solved, with complete explanations.

----------------

To see many other similar solved problems,  look into the lessons
    - Problem on a projectile moving vertically up and down
    - Problem on an arrow shot vertically upward
    - Problem on a ball thrown vertically up from the top of a tower
    - Problem on a toy rocket launched vertically up from a tall platform
in this site.